OpenStudy (anonymous):

Soybean is 14% protein; cornmeal s 7% protein, How many pounds of each should be mixed together to get 280-lb mixture that is 12% protein?

OpenStudy (radar):

Let x = lbs of Soybean Let y = lbs of cornmeal x+y=280 lbs .14x + .07y=.12(280)=33.6 x=280-y .14(280-y)+.07y=33.6 39.2-.14y+.07y=33.6 -.07y=-5.6 y=80 lbs x=200 lbs

OpenStudy (radar):

80 pounds cornmeal 200 pounds of soybean